OpenStudy (lina777):

1. csc^-1(csc(-pi/4))=-pi/4 TRUE OR FALSE and explain why 2. sec(sec^-1(sqrt3))=sqrt3 TRUE OR FALSE and explain why

OpenStudy (lina777):

@jtug6

Parth (parthkohli):

yes and yes

OpenStudy (aaronandyson):

^

OpenStudy (lina777):

Haha, but why?

OpenStudy (lina777):

I have to have an explanation for part of my answer

Parth (parthkohli):

that's how inverse functions work, really. csc(-pi/4) gives you the cosecant of -pi/4 (we don't really care what it is) then we're entering this value into the csc^(-1) function, which tells us the angle that has that cosecant. but we know the angle that has this cosecant: it's -pi/4. (there are other tidbits, like the inverse cosecant returns the angle in the range -pi/2 to pi/2, but fortunately here they happened to make the problem simple for us.)

OpenStudy (lina777):

Awesome thanks! @ParthKohli

Parth (parthkohli):

haha no problem, but can you come up with a similar explanation for the second?

OpenStudy (lina777):

Yes, it's is the same except for the inverse function is on the inside of the brackets... so the inverse is taken before the actual function is taken?

Parth (parthkohli):

yeah, but i hope that explanation above made sense.

OpenStudy (lina777):

Could I essentially say that they cancel each other out?

OpenStudy (lina777):

@ParthKohli

Parth (parthkohli):

yes, you can think of it that way. actually that's literally the definition of an inverse function. but be careful - \(\sin^{-1} (\sin x)\) is not always equal to \(x\) - likewise for other trigonometric functions too.

OpenStudy (lina777):

Okay, awesome! Haha. But why is that not always to x though? @ParthKohli

Parth (parthkohli):

alright, the reason is this: there are infinitely many angles that have the same sine, right? so if I asked you the inverse sine of 1, would you say pi/2? or 5pi/2? or -3pi/2? because all of these angles have their sin equal to 1. that is why inverse sine has been defined as the angle BETWEEN \(-\pi/2\) and \(\pi/2\) which has the sine equal to the input... this ensures that there is only one angle returned for a given sine! so could you tell me the value of \(\sin^{-1}(\sin (2\pi/3))\)?

OpenStudy (lina777):

Oh I see! So, it would have to be between that range, therefore, it would be pi/3 because that is the same value?

OpenStudy (lina777):

@ParthKohli

Parth (parthkohli):

Yes, amazing! So now you see that you don't get the same angle.

Parth (parthkohli):

Adding to this, can you guess the condition for \(\sin^{-1}(\sin(x))= x \)?

OpenStudy (lina777):

I'm sorry, but what do you mean by the condition? @ParthKohli

Parth (parthkohli):

Well, for what values of \(x\) does \(\sin^{-1}(\sin (x)) = x\) hold true? In other words, I'm just asking you to solve the equation.

OpenStudy (lina777):

Oh, I see. It would merely be the range of sin, right?

OpenStudy (lina777):

I also have another question: so for this study guide I am doing I am asked to solve problems such as 2cosx+5=4. That simplifies to cosx=-1/2. So, when I am evaluating x, would I say that the answer would be x=2/3pi +/- 2npi and x=4/3pi +/- 2npi?

Parth (parthkohli):

Range of sin?

OpenStudy (lina777):

-pi/2 to pi/2? or -1 to 1

OpenStudy (aaronandyson):

-1to 1

Parth (parthkohli):

You must have meant the range of arcsine.

OpenStudy (lina777):

oh yes!! haha sorry.

OpenStudy (lina777):

I do have a question though, for question 1. Wouldn't it be false because the range of csc is (-infinity, -1) U (1 to infinity) and the value is pi/4?

Parth (parthkohli):

But we're talking about inverse cosec here.

OpenStudy (lina777):

I'm sorry... so is there a different range for the inverse of arccsc?

OpenStudy (lina777):

I'm so sorry to be taking so much of your time @ParthKohli I really appreciate it!

Parth (parthkohli):

Yes, it's -pi/2 to pi/2 (excluding zero, of course).

Parth (parthkohli):

Inverse trig functions return an angle.

OpenStudy (lina777):

Oh great! Ha. I did not know that was a thing. So sorry, but one more thing: I also have another question: so for this study guide I am doing I am asked to solve problems such as 2cosx+5=4. That simplifies to cosx=-1/2. So, when I am evaluating x, would I say that the answer would be x=2/3pi +/- 2npi and x=4/3pi +/- 2npi?

OpenStudy (lina777):

So what I had put down would be the answer?

Parth (parthkohli):

OK, I'm dumb right now.

OpenStudy (lina777):

Wow. Thank you so very much. I can't tell you how much this has clarified my problems! I pride myself in math, but sometimes you just have those moments where you cannot figure things out (or second guess yourself haha). So seriously, thanks!

Parth (parthkohli):

The condensed form is \(2n \pi \pm \frac{2\pi}3\). I just confused myself thrice.

OpenStudy (lina777):

hahaha. But my answer will be marked as correct, correct?

Parth (parthkohli):

yeahhh I guess